If r = x
y
1
1 0 g
and A = 0 1 f
g f c
show that r^T Ar = 0 is the equation of the
circle in R^2 with centre (−g, −f) and radius square root(g^2 + f^2 − c).
1
Expert's answer
2020-06-01T18:28:08-0400

"r=\begin{pmatrix}
 x \\
 y\\1
\end{pmatrix}"

"r^T=\begin{pmatrix}
 x & y&1
\end{pmatrix}"

"A=\begin{pmatrix}
 1 &0&g \\
 0 & 1&f\\
g&f&c
\end{pmatrix}"

"r^TAr=\begin{pmatrix}
 x & y&1
\end{pmatrix}\begin{pmatrix}
 1 &0&g \\
 0 & 1&f\\
g&f&c
\end{pmatrix}\begin{pmatrix}
 x \\
 y\\1
\end{pmatrix}
=\\
=\begin{pmatrix}
 x+g&y+f&xg+yf+c
\end{pmatrix}\begin{pmatrix}
 x \\
 y\\1
\end{pmatrix}=\\
=(x^2+xg+y^2+yf+xg+yf+c)=\\
=(x^2+2xg+y^2+2yf+c)=0"

"x^2+2xg+g^2-g^2+y^2+2yf+f^2-f^2+c=0\\
(x+g)^2+(y+f)^2=g^2+f^2-c"

The circle: centre "(-g,-f)" , radius "R=\sqrt{g^2+f^2-c}"